OpenStudy (anonymous):

PLEASE HELP =) Use the quadratic formula to solve for x 3x^2 + 5x =3 Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas.

OpenStudy (anonymous):

You have the equation in the form Ax^2+BX+C=0, you can just plug your numbers into the formula go to town

OpenStudy (anonymous):

I've never learned this before though so i'm not sure what it's talking about ( or I don't remember either way)

OpenStudy (anonymous):

\[\frac{ -B \pm \sqrt{B ^{2}-4AC} }{ 2A }\]

OpenStudy (anonymous):

and I use my equation to replace the variables?

OpenStudy (anonymous):

what replaces what?

OpenStudy (anonymous):

-2 and 0.5?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

So as the person before said, you have to re-arrange the equation to look like this \(Ax^2+BX+C=0\) \(3x^2 + 5x =3 \rightarrow 3x^2 + 5x -3 =0 \) good so far? Next you use the quadratic formula \(\sf x= \dfrac{ -B \pm \sqrt{B ^{2}-4AC} }{ 2A }\) and you plug in the variables from your equation \(\underbrace{3}_Ax^2 + \underbrace{5}_Bx+ \underbrace{(-3)}_C =0\) so at the end it looks like this \(\sf x= \dfrac{ -(5) \pm \sqrt{5 ^{2}-4(3)(-3)} }{ 2*3 }\) then you solve for both values: \(\sf x= \dfrac{ -(5) + \sqrt{5 ^{2}-4(3)(-3)} }{ 2*3 }\) and \(\sf x= \dfrac{ -(5) - \sqrt{5 ^{2}-4(3)(-3)} }{ 2*3 }\)

OpenStudy (aaronq):

the computation is easiest if you do it in steps. First do the stuff under the squared root, then the addition or subtraction, then the division

OpenStudy (aaronq):

hope this helps

OpenStudy (anonymous):

thank you both! but yes that helped sooo much!I was just confused where the c in my equation was, at the top where it said +c why did it change to -c when we plugged in the numbers from the equation?

OpenStudy (anonymous):

so what I came out with so far is \[-5- \sqrt{\frac{ 61 }{ 6? }}\]

OpenStudy (anonymous):

@aaronq i'm still confused though because it doesn't give me the option in my answer box to type in fractions or an equation of any sort just one answer or commas between answers =o

OpenStudy (aaronq):

the -3 came from re-arranging the original equation: \(3x^2 + 5x =3 \rightarrow 3x^2 + 5x -3 =0\) subtract 3 from both sides: \(3x^2 + 5x -3 =3-3 \) \(3x^2 + 5x -3 =0 \) As for solving the equation, you are not doing it correctly. The addition/subtraction step needs to be performed BEFORE the division. The squared root sign is not part of the denominator. \(\sf x= \dfrac{ -(5) + \sqrt{5 ^{2}-4(3)(-3)} }{ 2*3 }\) \(\sf x= \dfrac{ -(5) + \sqrt{61} }{ 6 }\) \(\sf x= \dfrac{ -(5) + 7.81 }{ 6 }\) \(\sf x= \dfrac{ 2.81 }{ 6 }=0.4683\)

OpenStudy (anonymous):

ohhh ok I see where I went wrong =o thank you!

OpenStudy (aaronq):

no problem!